OpenStudy (anonymous):

Half - life Stronium - 90 has a half - life of 28 days.

OpenStudy (campbell_st):

do you have an equation that models the decay...?

OpenStudy (anonymous):

(a) A sample has a mass of 50 mg initially. Find a formula for the mass remaining after t days. (b) Find the mass remaining after 40 days. (c) How long does it take the sample to decay to a mass of 2mg? (d) Sketch the graph of the mass function.

OpenStudy (wolf1728):

I'll find an equation - gee what a clever message board this is :-)

OpenStudy (anonymous):

I just started this, so I'm kind of confused how to do it.

OpenStudy (campbell_st):

ok... so the model will be something like \[A = P \times e^{-kt}\]

OpenStudy (campbell_st):

so you know the A =0.5 when t = 28 and P = 1 P is the intial quantity.... A is the quantity after t time periods and k is the growth constant

OpenStudy (wolf1728):

Here's a good page (and calculator) concerning half-lives http://1728.org/halflife.htm has formulas too

OpenStudy (campbell_st):

so \[0.5 = 1 \times e^{-k \times 28}\] you'll need to find k, using logs

OpenStudy (campbell_st):

hope it makes sense

OpenStudy (anonymous):

Got ya

OpenStudy (campbell_st):

then when you know the value of k, you can then undertake the other parts of the question. But thats on the proviso that the decay is exponential... and uses the model above

OpenStudy (wolf1728):

for part b) Ending amount = beginning amount / 2^n where n = time / half-life Ending amount = 50mg / 2^(40/28) Ending amount = 50 mg / 2^1.4285714286 Ending amount = 50 mg / 2.6918003853 18.5749286142 mg

OpenStudy (anonymous):

Thanks wolf ^.^

OpenStudy (campbell_st):

well there you go, its a base 2 exponential equation rather than a base e equation...

OpenStudy (anonymous):

k = ln2/28

OpenStudy (wolf1728):

u r welcome iambatman

OpenStudy (anonymous):

-ln*

OpenStudy (wolf1728):

c) time to decay to 2 mg time = (half-life) * log(begng amt / ending amt) / log (2) time = 28 days * log(50/2) / log (2) time = 28 days * log (25) / log(2) time = 28 days * 4.6438561898 time = 130.0279733137 days or about 130 days

OpenStudy (anonymous):

Ah, thanks a lot!!

OpenStudy (anonymous):

Wolf for part b I don't understand why you're dividing by 2^...

OpenStudy (anonymous):

Thought it was just 50e^-0.025*40

OpenStudy (wolf1728):

Sorry for being away for a while. Does someone need a graph? (Okay now I'll work on your question).

OpenStudy (anonymous):

Lol awesome

OpenStudy (anonymous):

I get 18.394..

OpenStudy (anonymous):

A little different from yours haha

OpenStudy (wolf1728):

That's okay :-) As for the dividing by the two, basically when I wrote that calculator and those equations, I wanted something that could be stated in plain English and NOT ⌠ 0 --> ∞ as f(x) , etc (or whatever strange calculus symbols they use).

OpenStudy (wolf1728):

Is that for part b you are referring?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

I used the formula y = 50e^-0.025t

OpenStudy (wolf1728):

Where does that .025 come from?

OpenStudy (anonymous):

k = -ln2/28

OpenStudy (anonymous):

Got 130 days as well woo, nice one wolf!

OpenStudy (anonymous):

Thank you so much, especially for the graph!

OpenStudy (wolf1728):

okay I'm trying to redo part b) using your formula So that should be 50 * e ^ -(ln(2)/28) ? with ln(2) = 0.6931471806 and ln(2)/28 = 0.0247552564 so it would be e ^ -.0247552564 I get 0.9755486421 for that part

OpenStudy (wolf1728):

It's okay iambatman I'm still trying to see what the difference is between your answer and mine for part b)

OpenStudy (wolf1728):

Oh I see it should be 50e^-0.025*40 e^-0.9902102579 = 0.3714985723 times 50 equals 18.5749286142

OpenStudy (anonymous):

:p

OpenStudy (wolf1728):

So, I guess I was right?

OpenStudy (anonymous):

Oh no your number is still the same lol

OpenStudy (wolf1728):

But I calculated the number the second time using your formula (check my math),

OpenStudy (anonymous):

0.367879441

OpenStudy (anonymous):

You're using exactly 0.025 right?

OpenStudy (wolf1728):

Heck no - (I wanted to see how we got different numbers). So you said 0.367879441 whereas I calculated 0.3714985723

OpenStudy (anonymous):

hahaha maybe that's why

OpenStudy (anonymous):

I've been using exactly 0.025 XD

OpenStudy (anonymous):

getting -1

OpenStudy (wolf1728):

Heck I had a feeling I was correct because I rechecked my calculations with the web page calculator.

OpenStudy (wolf1728):

I get -0.9902102579

OpenStudy (anonymous):

Lol sorry about that, and did you use 0.025 and got that now?

OpenStudy (wolf1728):

No I used e^-0.9902102579 which equals 0.3714985723 then multiplying by 50 I got 18.5749286142

OpenStudy (wolf1728):

I have to get going be back in 15 minutes. By the way, please check out my web site www.1728.org it has some rather nice calculators

OpenStudy (anonymous):

Alright later wolf, and awesome, I will definitely check it out.

OpenStudy (anonymous):

Thanks again!